The value of ultimate period of oscillation Pu is 3minutesand the ultimate control gain Kcu is 2,select correct set of parameters for PID controller using Zeigler-nicholas

1
Expert's answer
2021-11-15T11:12:08-0500

Kp = 0.6Ku

τi = Pu/2

τd = Pu/8

Kp = Controller gain value that you should enter into the controller for good performance

Ku = “Ultimate” gain determined by increasing controller gain until self-sustaining oscillations are achieved

τi = Controller integral setting that you should enter into the controller for good performance (minutes per repeat)

τd = Controller derivative setting that you should enter into the controller for good performance (minutes)

Pu = “Ultimate” period of self-sustaining oscillations determined when the controller gain was set to Ku (minutes)

Kp = 0.6Ku = 0.6*2 = 1.2 minutes

τi = Pu/2 = 3/2 = 1.5 minutes per repeat

τd = Pu/8 = 3/8 = 0.375 minutes